2 kg of air at 5 bar, 80°C expands adiabatically in a closed system until its volume is doubled and its temperature become equal to that of the surrounding which is at 1 bar, 50°C. For the process, the maximum work is (if Cv= 0.718KJ/kg.k)
a) 23 KJ
b) 123 KJ
c) 223 KJ
d) 423 KJ
1
Expert's answer
2019-09-17T12:05:20-0400

The work in the adiabatic process is the change in internal energy, i.e."A = mc_v(T_1 - T_2) = 2\cdot0.718(80 - 50) = 43 (KJ),"

unfortunately it's out of listing.